Hello,

I need help using a system of equations to solve this problem:

Adult tickets for the school musical sold for $3.50 and student tickets sold for $2.50. Three hundred twenty-one tickets were sold altogether for $937.50. How many of each kind of tickets were sold?

Thank you

Answer by scott8148(5895)

a+s=321 ... s=321-a

350a+250s=93750 ... 350a+250(321-a)=93750